OpenStudy (anonymous):

PLEASE HELP! IM FAR BEHIND AND NEED HELP! MEDAL TO BEST ANSWER AND FANNED What number systems do the following numbers belong? [Natural Numbers, Whole Numbers, Integers, Rational Numbers, Irrational Numbers – There can be more than 1 answer.] 7. -6.7 8. √289 9. √5

OpenStudy (anonymous):

\[-6.7=-\frac{ 67 }{10 }=rational~ number.\] \[\sqrt{289}=17=?\](natural no.,whole no.,integer. all three) \[\sqrt{5}=irrational~ number.\]

OpenStudy (anonymous):

OMG THANK YOU SO MUCH!